OpenStudy (anonymous):

For the indicated functions f and g, find the functions f+g, f-g, fg, and f/g, and find their domains. f(x)=4x; g(x)=x+1

OpenStudy (anonymous):

f+g = 4x + x+1 = The domain are all real number f-g = 4x - x+1 = The domain are all real numbers fg = 4x * x-1 = The domain are all real numbers f/g = 4x / x+1 = I will let you figure this one out. What do you think the domain is?